Ravi has 3 hens that lay eggs. One hen lays 1 egg every day. The second hen lays 1 egg every day. The third hen lays 1 egg every 3 days. How many eggs would Ravi gather over a 6 days?
Found 2 solutions by JulietG, MathTherapy:
Answer by JulietG(1812)

Figure out how many eggs each hen lays in 6 days, and then add them together.
Hen A lays 1 per day. 1*6 = 6
Hen B lays 1 every 2 days. 1/2 * 6 = 3
Hen C lays 1 every 3 days. 1/3 * 6 = 2
.
Total = 6 + 3 + 2
=11 eggs in 6 days
(Well, crud, I read that as the second hen lays every OTHER day.)

Answer by MathTherapy(10806)

Ravi has 3 hens that lay eggs. One hen lays 1 egg every day. The second hen lays 1 egg every day. The third hen lays 1 egg every 3 days. How many eggs would Ravi gather over a 6 days?

1st hen: 6 eggs in 6 days
2nd hen: 6 eggs in 6 days
3rd hen: 2 eggs in 6 days
Total eggs: 14 eggs in 6 days
You can do the check!!
Send comments, “thank-yous,” and inquiries to “D” at MathMadEzy@aol.com.
Further help is available, online or in-person, for a fee, obviously